# A Three-Level Whole-Body Disturbance Rejection Control Framework for Dynamic Motions in Legged Robots

### v1

@@ II Preliminaries
We denote $h$ as the linear momentum and L{L} as the angular momentum around the robot’s center of mass (CoM). Denoting $c\in\mathbb{R}^{3}$ as the position of the CoM in the inertial frame, and $h=m\dot{c}$ denotes the linear momentum, where $m$ is the total mass of the robot. The centroidal momentum is linked to the joint configuration and its derivatives through the centroidal mapping
@@
The legged robot’s control system consists of several modules, including a two-level control (high-level control and low-level control), state estimation, and gait scheduler [22], as presented in Fig. 1. A reference trajectory can be generated for high-level control based on user input and state estimation. The gait scheduler determines the timing and sequence of the gait, coordinating the transition between the swing and stance phases of each leg. The high-level control module manages the position of the swing legs and the optimal GRFs for the stance legs, all in accordance with the user commands and gait timing. These components of the control architecture are detailed in [23, 24, 22].
@@
To address the uncertainties and noise in legged robots, we propose a three-level whole-body disturbance rejection control architecture, as illustrated in Fig. 2. We employ model predictive control (MPC) in high-level control to plan the legged robot’s centroidal momentum, joint positions, joint velocities, and GRFs based on nominal centroidal dynamics. The MH-ESO in mid-level control estimates disturbances and addresses noise in the legged system, while the MPC is used in mid-level control to replan the robot’s joint acceleration and GRFs, based on the reference trajectory from high-level control and accounting for uncertainties. Finally, whole-body control in low-level control compensates for system disturbances.
@@ III High-level Control
$\dot{x}_{c}=f_{c}(x_{c},u_{c})+d_{c}$ (8)
where $d_{c}\in\mathbb{R}^{12+n}$ accounts for the uncertainties in the centroidal dynamics. The $d_{c}$ represents the difference in uncertainties between the nominal centroidal dynamics and the centroidal dynamics with disturbances. The centroidal dynamics (8) represents the nominal centroidal dynamics when $d_{c}$ is a zero vector. The continuous dynamics (8) is discretized over intervals of the prediction horizon $[t,t+T]$ for the nonlinear trajectory optimization problem. We consider the total number of steps in the MPC to be $H$, and thus the duration of each discrete interval is $\delta_{t}=T/(H-1)$. Define $J_{c}(x)$ as the terminal quadratic cost around the reference state $x_{\rm{des}}$
$J_{c}(x)=(x-x_{\rm{des}}(T))^{\top}Q_{t}(x-x_{\rm{des}}(T))$ (9)
and $\varphi_{c}(x,u)$ as the quadratic cost around the reference state $x_{\rm{des}}$ and input $u_{\rm{des}}$
$\varphi_{c}(x,u)=$ $(x-x_{\rm{des}}(T))^{\top}Q_{s}(x-x_{\rm{des}}(T))$
$+(u-u_{\rm{des}}(T))^{\top}R(u-u_{\rm{des}}(T)).$ (10)
Let $x_{i,k}$ and $u_{i,k}$ denote the sequences of state and input variables in the MPC at the $k$-th solve, respectively, for $i=1,\cdots,H-1$. The nonlinear MPC problem can be formulated by defining and evaluating a cost function and constraints as [13, 10]
@@
where ${x}_{m,k}$ is the current measured state, $X=[x_{0,k}^{\top},\cdots,x_{H,k}^{\top}]^{\top}$ and $U=[u_{0,k}^{\top},\cdots,u_{H-1,k}^{\top}]^{\top}$ are the sequences of state and input variables respectively. The penalized cost function $l_{i}$ is derived from inequality constraints, which ensure that the swing leg remains above the ground and that the contact force lies within the friction cone. The constraint function $g_{i}$ represents the equality requirements, ensuring that the swing leg’s end does not experience a contact force and that the end of the stance leg has zero velocity. Note that the legged robot is a hybrid system, where the $f_{i}$ in the GRFs $F$ is either a zero vector or a nonzero vector, with this being managed by the gait scheduler.
The solution to the MPC problem (11) is denoted as $X=X_{\rm{ref}}:=[h_{\rm{ref}}^{\top},{L}_{\rm{ref}}^{\top},q_{\rm{ref}}^{\top}]^{\top}$ and $U=U_{\rm{ref}}:=[F_{\rm{ref}}^{\top},(\dot{q}^{j}_{\rm{ref}})^{\top}]^{\top}$, which gives the reference signals at the discrete time instances $t,t+\delta_{t},\cdots,t+T]^{\top}$. We can linearly interpolate the discrete reference signals to a continuous reference trajectory for the interval $[t,t+T]$. We use the nominal dynamics with $d_{c}$ being a zero vector in the MPC problem (11), since we prefer to generate the reference trajectory as if the system does not encounter the disturbances.
@@ IV Uncertainties Estimation
In this section, we introduce two disturbance estimators, ESO and MH-ESO, which is designed based on ESO, to estimate uncertainties in the legged system, as discussed in section IV-A. Both disturbance estimators can be adopted in our proposed T-WB-DRC framework. We also present the signal processing module in T-WB-DRC, as described in section IV-B.
@@ IV-A Disturbance Estimator
We will begin by designing the ESO and then integrate it into the moving horizon estimation (MHE) process to effectively manage noise.
@@ IV-A1 ESO design
We extend the state $x_{3}\in\mathbb{R}^{6+n}$. Let $x_{3}=d_{e}$ and $h_{x}(t)$ denote the rate of change of uncertainty, i.e., $h_{x}(t)=\dot{x}_{3}$, both $x_{3}$ and $h_{x}(t)$ are assumed to be unknown but bounded vector functions. The system (20) can be expressed in state-space form as
@@
where $\hat{u}=[\tau^{\top},F_{\rm{ref}}^{\top}]^{\top}$ with $F_{\rm{ref}}$ obtained by solving the MPC problem (11) and $\omega_{0}$ is tuning parameter to be determined. We use $F_{\rm{ref}}$ in $\hat{u}$ as the system input to estimate the system uncertainties. The discrepancy between $F$ and $F_{\rm{ref}}$ is estimated using the ESO.
@@
[…]
Note that the matrix $A$ is Hurwitz.
@@
Consider the error dynamics (30) under Assumption IV.1. Suppose both the disturbance $x_{3}$ and its derivative $h_{x}(t)$ are bounded, i.e., $x_{3}||\leq d_{b}$ and $h_{x}(t)||\leq h_{b}$, then the estimation error $\eta$ is bounded.
Proof: Let us consider the following control Lyapunov candidate function:
$V(\eta)=\eta^{\top}P\eta$ (39)
$+\frac{{{h_{x}^{\top}}{C_{2}}^{\top}P\eta+{\eta^{\top}}P{C_{2}}h_{x}}}{{{\omega_{0}}^{2}}}.$ (40)
It follows that
@@
From (39), we have
[…]
where
[…]
From (43), we have
$\dot{V}(\eta)+\lambda V(\eta)\leq\lambda{\delta_{V}}.$ (44)
Invoking the comparison lemma [27] for (44) yields
where $\eta_{0}$ is the values of $\eta$.
[…]
Thus, the estimation error $\eta$ is bounded.
By adopting the ESO, the estimated uncertainties $\hat{f}$ of legged robots are given by
@@ IV-A2 MH-ESO design
The ESO in (52) can be discretized using the backward Euler formula, yielding as
@@ V Robust Model Predictive Control
In this section, we employs a robust MPC based on centroidal dynamics with uncertainties to replan the GRFs reference trajectories for low-level control.
Due to the uncertainties are estimated in section IV based on whole-body dynamics, we first calculate the uncertainties in the centroidal dynamics using the uncertainties $d$ derived from the whole-body dynamics (1). Note that the $d_{c}$ represents the differences in uncertainties between the centroidal dynamics with disturbances and the nominal centroidal dynamics. The derivatives of centroidal momentum are obtained as [20]
@@
Thus, the uncertainties $d_{c}$ is
${d}_{c}=f_{h}(q,\dot{q},\ddot{q})-f_{h}(q_{\rm{ref}},\dot{q}_{\rm{ref}},\ddot{q}_{\rm{ref}}).$ (107)
It is noted that $q_{\rm{ref}}$, $\dot{q}_{\rm{ref}}$, $\ddot{q}_{\rm{ref}}$ are derived from the nominal MPC in high-level control.
Subsequently, we construct a robust MPC approach to determine the reference trajectories for the robot’s GRFs, while accommodating uncertainties. The continuous dynamics (7) are discretized over intervals of the prediction horizon $t,t+T_{0}]$ for the formulation of the nonlinear MPC trajectory optimization problem. Choosing the time horizon of the MPC problem as M1M-1, thus the discrete interval is set to $\delta_{t_{0}}=T_{0}/(M-1)$. Define
$J_{u}(x)=(x-X_{\rm{ref}}(t))^{\top}Q^{u}_{t}(x-X_{\rm{ref}}(t))$
and
$\varphi_{u}(x,u)=$ $(x-X_{\rm{ref}}(t))^{\top}Q^{u}_{s}(x-X_{\rm{ref}}(t))$
$+(u-U_{\rm{ref}}(t))^{\top}R(u-U_{\rm{ref}}(t)).$
where $X_{\rm{ref}}$ and $U_{\rm{ref}}$ are obtained by solving the nominal MPC problem in high-level control. Since the acceleration $\ddot{q}$ in (103) cannot be measured directly, we obtain the acceleration estimation $\ddot{\hat{q}}$ as
@@
with $\hat{f}_{\rm{process}}$ being given in (96). From (107), the uncertainty estimate $\hat{d}_{c}$ is given by
$\hat{d}_{c}=f_{h}(q,\dot{q},\ddot{\hat{q}})-f_{h}(q_{\rm{ref}},\dot{q}_{\rm{ref}},\ddot{q}_{\rm{ref}}).$ (108)
The nonlinear robust MPC trajectory optimization problem can be formulated as
@@
$x_{i+1,k}-x_{i,k}-f_{c}(x_{i,k},u_{i,k})\delta t_{0}=\hat{\textbf{{d}}_{\textbf{c}}}$,
@@ VII Simulation Verification
To evaluate the effectiveness and versatility of the T-WB-DRC, we conduct a series of simulation experiments on legged robots. First, we present the advantages of the MH-ESO over the ESO. Second, we compare the performance of the proposed T-WB-DRC with the WB-DRC introduced in [13] and the widely adopted standard whole-body control framework (standard WB-C) from [10], considering model uncertainties, external disturbances, and fault tolerance.
@@ VII-A Impact of Noise on ESO and MH-ESO
We conduct a simulation on the robot carrying a 10 kg load and standing in place to compare the Mean Squared Error (MSE) and the variance of the Frobenius norm of the estimated uncertainties $\hat{f}$ using MH-ESO and ESO. This comparison demonstrates the impact of noise on the estimated uncertainties using both MH-ESO and ESO while the robot remains stationary with the 10 kg load. The MSE and variance of the Frobenius norm of the estimated uncertainties are defined as

### v2

@@ Abstract
Note to Practitioners
This paper presents a practical control framework to significantly improve the robustness of legged robots against real-world uncertainties like unknown payloads, external pushes, and actuator faults. Its core is a novel three-level whole-body controller (T-WB-DRC) that uses a moving horizon estimator (MH-ESO) to accurately identify and compensate for disturbances in real-time. This dual-planning approach, which considers both ideal and disturbance-injected dynamics, outperforms previous methods. The framework’s effectiveness for enhancing stability under disturbance has been successfully validated through extensive simulations and physical experiments on a quadruped robot.
@@ II Preliminaries
We denote $h$ as the linear momentum and L{L} as the angular momentum around the robot’s center of mass (CoM). Let $c\in\mathbb{R}^{3}$ be the position of the CoM in the inertial frame, and $h=m\dot{c}$ be the linear momentum, where $m$ is the total mass of the robot. The centroidal momentum is linked to the joint configuration and its derivatives through the centroidal mapping
@@
The legged robot’s control system normally consists of several modules, including a two-level control (high-level control and low-level control), state estimation, and gait scheduler [22], as presented in Fig. 1. A reference trajectory can be generated by the high-level control based on the user input and the state estimation. The gait scheduler determines the timing and sequence of the gait, coordinating the transition between the swing and stance phases of each leg. The high-level control module manages to generate the trajectory for the swing legs and the optimal GRFs for the stance legs, all in accordance with the user commands and gait timing. These components of the control architecture are detailed in [23, 24, 22].
@@
To address the uncertainties and noise in legged robots, we propose a three-level whole-body disturbance rejection control architecture, as illustrated in Fig. 2. We employ model predictive control (MPC) in the high level to plan the legged robot’s centroidal momentum, joint positions, joint velocities, and GRFs based on nominal centroidal dynamics. The MH-ESO in middle level estimates disturbances and addresses the noise in the legged system, while the MPC is then used to replan the robot’s joint acceleration and GRFs, based on the reference trajectory from high-level control. Finally, the whole-body control in the low level compensates for system disturbances.
@@ III High-level Control
$\dot{x}_{c}=f_{c}(x_{c},u_{c})$ (8)
The continuous dynamics (8) is discretized over intervals of the prediction horizon $[t,t+T]$ for the nonlinear trajectory optimization problem. We consider the total number of steps in the MPC to be $H$, and thus the duration of each discrete interval is $\delta_{t}=T/(H-1)$. Define $J_{c}(x)$ as the terminal quadratic cost around the reference state $x_{\rm{des}}$
$J_{c}(x)=(x-x_{\rm{des}}(t+T))^{\top}Q_{t}(x-x_{\rm{des}}(t+T))$ (9)
and $\varphi_{c,i}(x,u)$ as the quadratic cost around the reference state $x_{\rm{des}}(t)$ and input $u_{\rm{des}}(t)$
$\varphi_{c,i}(x,u)=$ $(x-x_{\rm{des}}(t+i\delta_{t}))^{\top}Q_{s}(x-x_{\rm{des}}(t+i\delta_{t}))$
$+(u-u_{\rm{des}}(t+i\delta_{t}))^{\top}R_{s}(u-u_{\rm{des}}(t+i\delta_{t}))$
for $i=1,\cdots,H-1$. Let $x_{i,k}$ and $u_{i,k}$ denote the sequences of state and input variables in the MPC at the $k$-th solve, respectively. The nonlinear MPC problem can be formulated by defining and evaluating a cost function and constraints as [13, 10]
@@
where ${x}_{m,k}$ is the current measured state, $X=[x_{0,k}^{\top},\cdots,x_{H,k}^{\top}]^{\top}$ and $U=[u_{0,k}^{\top},\cdots,u_{H-1,k}^{\top}]^{\top}$ are the sequences of state and input variables respectively. The penalized cost function $l_{i}$ is derived from inequality constraints, which ensure that the swing leg remains above the ground and that the contact force lies within the friction cone. The constraint function $g_{i}$ represents the equality requirements, ensuring that the swing leg’s end does not experience a contact force and that the end of the stance leg has zero velocity. Note that the legged robot is a hybrid system, when managed by the gait scheduler, $f_{i}$ in the GRFs $F$ is either a zero vector or a nonzero vector.
The centroidal dynamics (8) represents the nominal centroidal dynamics. We use the nominal dynamics in high-level control, as we prefer to generate the reference trajectory under the assumption that the system does not encounter disturbances. The solution to the MPC problem (10) is denoted as $X=X_{\rm{ref}}:=[h_{\rm{ref}}^{\top},{L}_{\rm{ref}}^{\top},q_{\rm{ref}}^{\top}]^{\top}$ and $U=U_{\rm{ref}}:=[F_{\rm{ref}}^{\top},(\dot{q}^{j}_{\rm{ref}})^{\top}]^{\top}$, which gives the reference signals at the discrete time instances $t,t+\delta_{t},\cdots,t+T]^{\top}$. We can linearly interpolate the discrete reference signals to a continuous reference trajectory for the interval $[t,t+T]$.
@@ IV Uncertainties Estimation
In this section, we introduce two variants of disturbance estimators, namely ESO and MH-ESO which is designed based on ESO, to estimate uncertainties in the legged system. The disturbance estimators are discussed in Section IV-A. Both disturbance estimators can be adopted in our proposed T-WB-DRC framework. We also present the signal processing module in T-WB-DRC in Section IV-B.
@@ IV-A Disturbance Estimator
The influence of measurement noise on disturbance estimators is always present in legged robots. For almost all disturbance estimators in legged robots, the measurement noise is not directly considered in the estimator design; instead, it is handled in the signal processing module using an MAF filter or low-pass filter. In this section, we will start by designing the ESO without considering the influence of measurement noise and then integrate it into the moving horizon estimation (MHE) process to effectively manage the noise in the estimator design.
@@ IV-A1 ESO design
We extend the state with $x_{3}\in\mathbb{R}^{6+n}$. Let $x_{3}=d_{e}$ and the time-varying signal $h_{x}$ denote the rate of change of uncertainty $d_{e}$, i.e., $\dot{x}_{3}=h_{x}(t)$. $x_{3}$ and $h_{x}(t)$ are assumed to be unknown but bounded vector functions. The system (19) can be expressed in the state-space form as
@@
where $\hat{u}=[\tau^{\top},F_{\rm{ref}}^{\top}]^{\top}$ with $F_{\rm{ref}}$ being the partial solution of the MPC problem (10) and $\omega_{0}$ is tuning parameter to be determined. We use $F_{\rm{ref}}$ instead of $F$ in (27), because we would like to estimate the discrepancy between $F$ and $F_{\rm{ref}}$ using the ESO.
@@
where $\tilde{u}$ is an unknown but a bounded vector, due to that $F_{\rm{ref}}$ is, $g_{e}(x_{1})$ is a bounded matrix, and
@@
Note that the matrix $A$ is Hurwitz. The following assumption is considered.
@@
The proof of Theorem IV.1 is provided in Appendix.
Note that $d_{e}=D^{-1}d$ in (19) is estimated by $\hat{x}_{3}$ in ESO. The estimated $\hat{f}$ for the uncertainties $d$ in (1) are given by
@@ IV-A2 MH-ESO design
The ESO in (43) can be discretized using the backward Euler formula, yielding
@@ V Robust Model Predictive Control
In this section, we aim to track the reference trajectory $\dot{X}_{\rm{ref}}:=[\dot{h}_{\rm{ref}}^{\top},\dot{L}_{\rm{ref}}^{\top},\dot{q}_{\rm{ref}}^{\top}]^{\top}$ generated by high-level control, by adjusting the GRFs of the centroidal dynamics using robust MPC.
We first calculate the change in the momentum derivatives of the legged system when uncertainties are present, compared to when the system is free of uncertainties, as follows:
[…]
The derivatives of centroidal momentum are obtained as [20]
@@
Since the acceleration $\ddot{q}$ in (101) cannot be measured directly, we obtain the acceleration estimation $\ddot{\hat{q}}$ as
@@
with $\hat{f}_{\rm{process}}$ being given in (87). From (94), the estimated difference in momentum derivatives, $\hat{d}_{c}$, is given by
[…]
Subsequently, we construct a robust MPC approach to determine the reference trajectories for the robot’s GRFs, while accommodating uncertainties. $d_{c}$ is considered for the compensation in the centroidal dynamics,
$\dot{x}_{c}=f(x,u)+d_{c},$ (109)
to track the reference trajectory $\dot{X}_{\rm{ref}}$. We use the estimation $\hat{d}_{c}$ to replace $d_{c}$ in the compensated centroidal dynamics for the robust MPC. The continuous dynamics (7) are discretized over intervals of the prediction horizon $t,t+T_{0}]$ for the formulation of the nonlinear MPC trajectory optimization problem. Choosing the time horizon of the MPC problem as M1M-1, thus the discrete interval is set to $\delta_{t_{0}}=T_{0}/(M-1)$. Define
$J_{u}(x)=(x-X_{\rm{ref}}(t+T_{0}))^{\top}Q^{u}_{t}(x-X_{\rm{ref}}(t+T_{0}))$
and
$\varphi_{u,i}(x,u)=$ $(x-X_{\rm{ref}}(t+i\delta_{t_{0}}))^{\top}Q^{u}_{s}(x-X_{\rm{ref}}(t+i\delta_{t_{0}}))$
$+(u-U_{\rm{ref}}(t+i\delta_{t_{0}}))^{\top}R(u-U_{\rm{ref}}(t+i\delta_{t_{0}})).$
where $X_{\rm{ref}}$ and $U_{\rm{ref}}$ are obtained by solving the nominal MPC problem in high-level control. The nonlinear robust MPC trajectory optimization problem can be formulated as
@@
$x_{i+1,k}-x_{i,k}-f_{c}(x_{i,k},u_{i,k})\delta t_{0}={\hat{\textbf{{d}}_{\textbf{c}}}\delta t_{0}},$
@@ VII Simulation Verification
To evaluate the effectiveness and versatility of the T-WB-DRC, we conduct a series of simulation experiments on legged robots. First, we present the advantages of the MH-ESO over the ESO. Second, we compare the performance of the proposed T-WB-DRC with the WB-DRC introduced in [13] and the widely adopted standard whole-body control framework (standard WB-C) from [10], while model uncertainties, external disturbances, and fault tolerance are considered.
@@ VII-A Impact of Noise on ESO and MH-ESO
We conduct a simulation on the robot carrying a 10 kg load and standing in place and compare the Mean Squared Error (MSE) and the variance of the Frobenius norm of the estimated uncertainties $\hat{f}$ using MH-ESO and ESO. This comparison demonstrates the impact of noise on the estimated uncertainties using both MH-ESO and ESO while the robot remains stationary with the 10 kg load. The MSE and variance of the Frobenius norm of the estimated uncertainties are defined as
@@ IX Conclusion
Proof of Theorem IV.1: Let us consider the following control Lyapunov candidate function:
$V(\eta)=\eta^{\top}P\eta$ (126)
$+\frac{{{h_{x}^{\top}}{C_{2}}^{\top}P\eta+{\eta^{\top}}P{C_{2}}h_{x}}}{{{\omega_{0}}^{2}}}.$ (127)
It follows that
[…]
From (126), we have
[…]
where
[…]
From (130), we have
$\dot{V}(\eta)+\lambda V(\eta)\leq\lambda{\delta_{V}}.$ (131)
Invoking the comparison lemma [29] for (131) yields
where $\eta_{0}$ is the values of $\eta$.
[…]
Thus, the estimation error $\eta$ is bounded.